Ana is 24‐year old when her son is born. In how many years

will Ana be thrice as old as her son?


1
Expert's answer
2021-09-28T08:38:33-0400

In 12 years.

Ana will be 36-year old.

"\frac{36}{3}=12"

or

"24 + x = x \times 3 \\

24 = 3x-x \\

24 = 2 x \\

x = \frac{24}{2} = 12"